OpenStudy (anonymous):

solve sec x csc x- 2 csc x = 0

OpenStudy (anonymous):

cosecx(secx-2)=0 sinx=0 or cosecx=0.5

OpenStudy (callisto):

sec x csc x- 2 csc x = 0 csc x (sec x- 2) = 0 csc x =0 or sec x -2 =0 x = 0 or x=180 or x= 360 or x=60 or x=300

OpenStudy (anonymous):

Should be cos x = 0.5 @Sarkar

OpenStudy (anonymous):

yes sorry for the error....